Calculation of individual costs and WACC Dillon Labs has asked its financial manager to measure the cost of each specific typPlease help with part c & d! Show your work ! Thanks
Answer #1

c)
Growth rate can be calculared by using the following excel formula:
=RATE(nper,pmt,pv,fv)
=RATE(10,0,-2.50,4.48)
= 6.01%

D1 = $4.48 * (1 + 6.01%) = $4.749

Cost of retained earnings = D1 / P0 + g
= ($4.749 / $60) + 6.01%
= 13.92%

Cost of retained earnings = 13.92%

Cost of new common stock = ($4.749 / ($60 - $2.50 - $2.50)) + 6.01% = 14.64%

Cost of new common stock = 14.64%

Average = (14.64% + 13.92%) / 2 = 14.28%

d)
WACC = (weight of debt * cost of debt) + (weight of equity * cost of equity) + (weight of preferred stock * cost of preferred stock)
= (40% * 6.16%) + (45% * 14.28%) + (15% * 8.62%)
= 10.18%

WACC = 10.18%